### v1

@@ 2 Related Work
The literature is rich with attempts to improve on the qualities and capabilities of the traditional Transformer architecture design [22],[23],[24],[25]. Guo et al. introduced Star Transformer [26], combining band attention and global attention. This formulation of the Transformer has a global node on which a band attention of width 3 is applied. Also, a shared global node connects a pair of non-adjacent nodes, while adjacent nodes are connected to each other. Beltagy et al. introduced Longformer[27], which also uses a combination of band attention and internal global-node attention. Classification tokens are selected as global nodes. The architecture substitutes the band attention heads in the upper layers with dilated window attention, thus increasing the receptive field without increasing computation. Kitaev et al. introduced Reformer [28] as a modified Transformer that employs locality-sensitive hashing (LSH). The LSH is used to select the key and value pairs for each query, therefore allowing each token to attend to tokens that exist in the same hashing bucket. BigBird architecture by Zaheer et al.[29] utilizes random attention to approximate full attention with a sparse encoder and sparse decoder, and it was shown by the analysis that this design can simulate any Turing Machine, explaining the capability of such architecture. Katharopoulos et al. proposed the Linear Transformer [30] with feature maps that target an approximation of the full Softmax Dot Product Attention and showed comparable performance in empirical tests.Wang et al. introduced Linformer [16], showing an approximation to the Attention mechanism by a low-rank matrix, thus lowering the computational requirement while maintaining comparable performance.Choromanski et al. proposed Performer [31], which uses random feature maps as an approximate to the Traditional Attention function. Wang et al. introduced the Cascade Transformer [32] By using a sliding window attention, the window size is exponentially increased when increasing the number of layers, leading to a reduction in complexity. Li et al. introduced the LogSparse Transformer [33] that facilitates long-term dependency on time series analysis by using Eponym attention. Qiu et al. introduced BlockBERT [34], which uses block-wise attention to split the input sequence into non-overlapping blocks. Tay et al. introduced the sparse Sinkhorn attention [35]. This mechanism is essentially block-wise attention, but the keys are sorted block-wise, therefore learning the permutations. Dai et al. proposed the Transformer-XL[36]. This design uses a recurrence between the windows that is segment-based. by storing the representations of the previous window and storing them in first-in, first-out memory (FIFO). After this step, the Transformer-XL applies attention to the sorted representations that have been stored in memory. Clustered Attention, proposed by Vyas et al. [37] clusters the quires, then calculates the attention distributions for cluster centroids.Zhang et al. proposed PoolingFormer [38], which utilizes a two-level attention, a sliding window attention, and a compressed memory attention. The compressed memory module is used after first applying the sliding window attention, then applying a compressed memory module for the purpose of increasing the receptive field. Liu et al. proposed Memory Compressed Attention (MCA) [39], which complements local attention with strided convolution, thus reducing the number of keys and values. This allows the architecture to process much longer sequences compared to traditional Transformers.Xiong et al. used the Nyström method to modify Transformer with the introduction of Nyströmformer [20]. This design selects landmark nodes by the process of strided average pooling and then processes these selected queries and keys with an approximation to attention by the Nyström method. Funnel Transformer [40] was proposed by Dai et al. by employing a funnel-like encoder that has a gradual reduction of the hidden sequence length using pooling along the sequence dimension; the proper length is then restored with an up-sampling process.Max-ViT [19] was introduced by Tu et al., which repeats the basic building block over multiple stages. The basic block consists of two aspects: blocked local attention and dilated global attention. Ho et al. proposed the Axial Transformer [41]. This architecture computes a sequence of attention functions with each one applied along a single axis of the input, reducing the computational cost. Swin Transformer [10] is an architecture proposed by Liu et al., and this design reduced the cost by splitting the image input into non-overlapping patches. These patches are then embedded as tokens for processing by Attention. FNets [17] was introduced by Lee-Thorp et al., and it proposes an attention-free Transformer architecture that substitutes the Softmax Dot Product function used in each The Fourier sublayer applies a 2D DFT to the embedded input in two steps: one 1D DFT along the sequence dimension and another 1D DFT along the hidden dimension. gMLP [42] was introduced by Liu et al., and this architecture is comprised of a series of blocks that are homogeneous in size and width. Each block layout is highly reminiscent of inverted bottlenecks. Another feature of this architecture compared to traditional Transformers is that it does not require position embeddings. Local-ViT [18] was introduced by Li et al. This architecture incorporates 2D depth-wise convolutions instead of the feed-forward network as in ViT. This design choice was inspired by the inverted residuals of MobileNets. Synthesizer [43] was proposed by Tay et al. as an architecture that learns synthetic attention weights and does not rely on interactions between tokens. The results showed competitive performance in relation to other linear transformer designs.Transformer iN Transformer (TNT) [44] was introduced by Han et al. This design treats the input images in a similar manner to a paragraph of text and divides them into several patches as “visual sentences” and then further divides them into sub-patches as “visual words.” With this hierarchical division, the architecture is divided into conventional transformer blocks for extracting features and attentions on the visual sentence level, and then a sub-transformer is introduced in order to extract the features of smaller visual words.
@@ 3.2 Proposed Architecture
(a) Multi-Head Self Attention
@@ 4 Experiments and Results
Dataset
Test accuracy (ViT) (%)
Test accuracy (MLP-Mixer)(%)
Test accuracy (LocalViT)(%)
Test accuracy (Ninformer) (%) (Ours)
MNIST
97.12
97.73
97.79
98.61
CIFAR-10
65.74
70.12
77.71
81.59
CIFAR-100
34.87
39.16
41.61
53.78

### v2

@@ 2 Related Work
The literature is rich with attempts to improve on the qualities and capabilities of the traditional Transformer architecture design [22],[23],[24],[25]. Guo et al. introduced Star Transformer [26], combining band attention and global attention. This formulation of the Transformer has a global node on which a band attention of width 3 is applied. Also, a shared global node connects a pair of non-adjacent nodes, while adjacent nodes are connected to each other. Beltagy et al. introduced Longformer[27], which also uses a combination of band attention and internal global-node attention. Classification tokens are selected as global nodes. The architecture substitutes the band attention heads in the upper layers with dilated window attention, thus increasing the receptive field without increasing computation. Kitaev et al. introduced Reformer [28] as a modified Transformer that employs locality-sensitive hashing (LSH). The LSH is used to select the key and value pairs for each query, therefore allowing each token to attend to tokens that exist in the same hashing bucket. BigBird architecture by Zaheer et al.[29] utilizes random attention to approximate full attention with a sparse encoder and sparse decoder, and it was shown by the analysis that this design can simulate any Turing Machine, explaining the capability of such architecture. Katharopoulos et al. proposed the Linear Transformer [30] with feature maps that target an approximation of the full Softmax Dot Product Attention and showed comparable performance in empirical tests.Wang et al. introduced Linformer [16], showing an approximation to the Attention mechanism by a low-rank matrix, thus lowering the computational requirement while maintaining comparable performance.Choromanski et al. proposed Performer [31], which uses random feature maps as an approximate to the Traditional Attention function. Wang et al. introduced the Cascade Transformer [32] By using a sliding window attention, the window size is exponentially increased when increasing the number of layers, leading to a reduction in complexity. Li et al. introduced the LogSparse Transformer [33] that facilitates long-term dependency on time series analysis by using Eponym attention. Qiu et al. introduced BlockBERT [34], which uses block-wise attention to split the input sequence into non-overlapping blocks. Tay et al. introduced the sparse Sinkhorn attention [35]. This mechanism is essentially block-wise attention, but the keys are sorted block-wise, therefore learning the permutations. Dai et al. proposed the Transformer-XL[36]. This design uses a recurrence between the windows that is segment-based. by storing the representations of the previous window and storing them in first-in, first-out memory (FIFO). After this step, the Transformer-XL applies attention to the sorted representations that have been stored in memory. Clustered Attention, proposed by Vyas et al. [37] clusters the quires, then calculates the attention distributions for cluster centroids.Zhang et al. proposed PoolingFormer [38], which utilizes a two-level attention, a sliding window attention, and a compressed memory attention. The compressed memory module is used after first applying the sliding window attention, then applying a compressed memory module for the purpose of increasing the receptive field. Liu et al. proposed Memory Compressed Attention (MCA) [39], which complements local attention with strided convolution, thus reducing the number of keys and values. This allows the architecture to process much longer sequences compared to traditional Transformers.Xiong et al. used the Nyström method to modify Transformer with the introduction of Nyströmformer [20]. This design selects landmark nodes by the process of strided average pooling and then processes these selected queries and keys with an approximation to attention by the Nyström method. Funnel Transformer [40] was proposed by Dai et al. by employing a funnel-like encoder that has a gradual reduction of the hidden sequence length using pooling along the sequence dimension; the proper length is then restored with an up-sampling process.Max-ViT [19] was introduced by Tu et al., which repeats the basic building block over multiple stages. The basic block consists of two aspects: blocked local attention and dilated global attention. Ho et al. proposed the Axial Transformer [41]. This architecture computes a sequence of attention functions with each one applied along a single axis of the input, reducing the computational cost. Swin Transformer [10] is an architecture proposed by Liu et al., and this design reduced the cost by splitting the image input into non-overlapping patches. These patches are then embedded as tokens for processing by Attention. FNets [17] was introduced by Lee-Thorp et al., and it proposes an attention-free Transformer architecture that substitutes the Softmax Dot Product function used in each The Fourier sublayer applies a 2D DFT to the embedded input in two steps: one 1D DFT along the sequence dimension and another 1D DFT along the hidden dimension. gMLP [42] was introduced by Liu et al., and this architecture is comprised of a series of blocks that are homogeneous in size and width. Each block layout is highly reminiscent of inverted bottlenecks. Another feature of this architecture compared to traditional Transformers is that it does not require position embeddings. Local-ViT [18] was introduced by Li et al. This architecture incorporates 2D depth-wise convolutions instead of the feed-forward network as in ViT. This design choice was inspired by the inverted residuals of MobileNets. Synthesizer [43] was proposed by Tay et al. as an architecture that learns synthetic attention weights and does not rely on interactions between tokens. The results showed competitive performance in relation to other linear transformer designs. Transformer iN Transformer (TNT) [44] was introduced by Han et al. This design treats the input images in a similar manner to a paragraph of text and divides them into several patches as “visual sentences” and then further divides them into sub-patches as “visual words.” With this hierarchical division, the architecture is divided into conventional transformer blocks for extracting features and attentions on the visual sentence level, and then a sub-transformer is introduced in order to extract the features of smaller visual words. De et al. proposed Hawk and Griffin models [45]; these are hybrid models combining gated linear recurrences and local attention with good extrapolation capabilities.
@@ 3.2 Proposed Architecture
(b) Multi-Head Self Attention
@@ 4 Experiments and Results
Dataset Test accuracy (ViT) (%) Test accuracy (MLP-Mixer)(%) Test accuracy (LocalViT)(%) Test accuracy (Ninformer) (%) (Ours)
MNIST 97.12 97.73 97.79 98.61
CIFAR-10 65.74 70.12 77.71 81.59
CIFAR-100 34.87 39.16 41.61 53.78